# TNO, The Hague, The Netherlands

### v1

@@ 2.5 QAOA
There exists an algorithm that translates quantum annealing to the gate-based quantum computer: QAOA, Quantum Approximate Optimisation Algorithm farhi2014quantum , which is a type of variational quantum eigensolver (VQE) algorithm. The VQE is a quantum algorithm that is used often for quantum chemistry, quantum simulations and optimisation problems. It is a hybrid algorithm that uses both classical computers and quantum computers to find the ground state of a given physical system. In QAOA, the adiabatic evolution is approximated by a discretised function that yields an approximation of the ground state of the desired Hamiltonian. The goal of the QAOA algorithm is to find the lowest possible upper bound for the ground state. Here too, we start with an equal superposition over all solutions |$\ket{s}$. The algorithm then alternately applies the following two operations, operators, to this state:
$U(H_{0},\beta)=e^{-i\beta H_{0}}\textrm{ and }U(H_{1},\gamma)=e^{-i\gamma H_{1}},$
@@
|$\ket{\gamma,\beta}=U(H_{0},\beta_{p})U(H_{1},\gamma_{p})…U(H_{0},\beta_{1})U(H_{1},\gamma_{1})\ket{s}.$
Using a QC, the expected value $F_{p}=\bra{\gamma,\beta}H_{1}\ket{\gamma,\beta}$ can then be calculated, which gives an upper limit for the ground state. If $p\rightarrow\infty$ and given the right choice of angels $\beta_{p}$ and $\gamma_{p}$, this approximation will converge to the optimal solution - the exact ground state of the underlying problem. However, finding the right angles is not trivial and if p must be large for a good approximation, the effectiveness of the algorithm is still an uncertain factor. The QAOA has more freedom than QA in this regard. Another choice for the so-called mixing operator $U(H_{0},\beta)$, is the Quantum Alternating Operator Ansatz (Hadfield et al. 2019), which can limit the search space, by, for example, setting hard constraints. The QAOA is an example of a variational or hybrid algorithm. Classical parameters are optimised to create a quantum circuit that solves a problem. This class of algorithms are seen to possess the first practical application to gated quantum computers. After all, it only needs a limited number of qubits and the depth of the circuit (the time factor) is also limited if p is not chosen too large. In addition, variational algorithms are fairly resistant to noise, the limited stability of the qubits.
@@ 3.1 Methodology
For this work we created an overview that is as complete as possible until 2023. We used Google Scholar and Scopus as source and used the index terms ‘quantum algorithm’ and the category names of the logistic problems. Note that there exists also a lot of literature on ‘quantum inspired’ algorithms like quantum genetic algorithms xihuai2007application , quantum evolutionary algorithm xu2009using and regular simulated annealing, not using the QUBO formulation, such as chibeles2012simulated ; mousavi2013hybrid . The first quantum annealing work we found is from 2013 and 2015, with the works from chooinvestigating ; crispin2013quantum ; venturelli2015quantum . The main flow of research starts around 2020, as shown in Fig. 4.
@@ 3.2.1 VRP and TSP quantum solutions
One of the earlier works is by Crispin et al. crispin2013quantum . They propose an approach to solve the Capacitated Vehicle Routing Problem (CVRP) using a quantum annealing algorithm. CVRP is a variant of the VRP where vehicles have limited capacity. The paper discusses the design of a spin encoding scheme for CVRP and an empirical approach for tuning parameters in the quantum annealing algorithm. QA is introduced as a metaheuristic that utilises quantum tunnelling in the annealing process. Experiments are conducted on CVRP benchmark instances, and the results are compared with a simulated annealing algorithm. The quantum annealing algorithm is found to match or outperform simulated annealing in terms of success rate for various instances. The paper analyses the quantum annealing algorithm’s performance and discusses the impact of parameters such as population size, temperature, and acceptance ratio.
@@
Srinivasan et al. srinivasan2018efficient , introduce a quantum algorithm to solve the TSP. It utilises the quantum phase estimation technique jiang2021survey , encoding distances between cities as phases. Unitary operators are constructed based on these phases, and the phase estimation algorithm is applied to estimate eigenvalues representing total distances for all possible routes. Quantum search algorithms are then employed to find the minimum distance and corresponding route, providing a quadratic speedup over classical brute force methods. The algorithm is illustrated with an example of TSP involving four cities, and simulations are conducted using IBM’s quantum simulator.
@@
The work by Feld et al. feld2019hybrid also investigates the CVRP using D-Wave’s quantum annealer. They propose a quantum-classical hybrid solution, incorporating a 2-Phase-Heuristic that divides CVRP into clustering and routing phases. Results indicate that the hybrid method can compete with classical construction and 2-phase heuristics in terms of solution quality. However, challenges arise, particularly in finding the ‘best known solution’ for certain datasets, emphasising the need for larger quantum hardware. The study anticipates future improvements in D-Wave’s technology, expecting increased qubit connectivity and quantity.
@@
Borowski et al. borowski2020new also delve into (C)VRP, introducing four quantum annealing-based algorithms. Their Full QUBO Solver (FQS), Average Partition Solver (AVS), DBSCAN Solver (DBSS), and Solution Partitioning Solver (SPS) showcase promising results. The study employs D-Wave’s Leap framework, comparing quantum algorithms with classical metaheuristics. The experiments use established benchmark datasets and custom scenarios with realistic road networks. The authors outline future research directions, including extending quantum formulations to address more realistic variants of VRP.
@@
Weinberg et al. weinberg2023supply focus on practical quantum algorithm use, considering the limitations of Noisy Intermediate-Scale Quantum (NISQ) hardware. They adopt a hybrid workflow for solving the CVRP, breaking down problems into smaller instances suitable for quantum annealing. Simulated annealing and the D-Wave Hybrid solver are used as quantum substitutes. While not claiming clear advantages over classical methods, they emphasise the sensible application of quantum algorithms to specific bottlenecks, anticipating future advancements.
@@
This study by Sinno et al. sinno2023performance focus on the performance of commercial QA solvers for the CVRP. The research challenges assumptions made in theoretical studies and simulations on classic hardware, emphasising the need for empirical measurements on real quantum platforms to assess commercial capabilities accurately. The study evaluates the quality of solutions provided by the D-Wave CQM solver for CVRP, considering problem size and complexity. The study suggests that, more than problem size, model complexity significantly affects solution quality, highlighting the importance of minimising constraint density in practical applications.
@@
Sales et al. sales2023adiabatic address the VRP in last-mile logistics, proposing a hybrid solution combining quantum annealing and classical techniques. Their analysis, using Amazon Braket and D-Wave’s QA, demonstrates the superiority of the hybrid algorithm over full Quantum Annealing, suggesting potential advantages in scenarios with lower data “clusterability.” Despite current challenges, the study identifies a cost-effective relationship between running quantum algorithms and result quality, indicating quantum computing’s theoretical advantage in solving constrained clustering problems.
@@
The thesis by Palmieri palmieriquantum explores the application of quantum optimisation techniques to address the (C)VRP. Here, a heuristic two-phase approach is explored. Nodes are first clustered into disjoint groups, and routing is performed individually within each cluster. Various classical methodologies for the clustering phase are investigated. An alternative approach models the clustering phase as a Modularity Maximization problem, which is tackled using quantum algorithms. The routing phase is modelled as a Travelling Salesman Problem (TSP), and subtour elimination constraints are adaptively added using the Dantzig-Fulkerson-Johnson (DFJ) strategy. This adaptive approach reduces the number of qubits, making the models suitable for both gate-based and annealing-based quantum devices. Various quantum algorithms, such as QA, quantum GAVA (Graver Augmented Multi-seed Algorithm), QAOA and VQE are investigated for quantum optimisation. Gurobi is used as a classical benchmarking method. QA and quantum GAMA outperfom the VQE approaches here.
@@
Until here, all works use Quantum Annealing as their solver for the VRP. Rana et al. rana2022quantum explore employee transport route and agri-tech logistics optimisation using both a quantum gate-based variational algorithm and QA. Here again, quantum hardware limitations lead to problem simplifications, using problem-specific decomposition techniques, such as distance-based clustering. The study utilises real-world data and anticipates extending quantum methods to more complex scenarios, acknowledging ongoing improvements in the D-Wave’s system for addressing larger, intricate logistics optimisation problems.
@@
Azad et al. azad2022solving explore the application of the QAOA for the VRP within Intelligent Transportation Systems (ITS). The paper provides an Ising formulation for the VRP and its variants, adapting them for quantum processing. Despite the limitations of NISQ devices, having 15 qubits, QAOA holds promise for solving combinatorial optimisation problems like VRP. However, they conclude that there is a need for a multi-faceted approach to enhance its performance. Working on aspects such as initialisation techniques, optimising parameter values, exploring different mixer Hamiltonians, and addressing noise-related challenges are essential for realising the full potential of QAOA in practical applications.
@@
The proposed Indirect QAOA (IQAOA) by Bourreau et al. bourreau2023indirect , introduces a novel approach, targeting the Hamiltonian of a set of string vectors. It utilises a Quantum Alternating Operator Ansatz with a parameterised family of unitary operators, allowing for efficient modelling of the Hamiltonian. This algorithm addresses the limitations of standard QAOA by incorporating a classical meta-optimisation loop and estimating the average cost of each string vector. The IQAOA approach includes the creation of a quantum circuit with a limited number of gates suitable for current noisy quantum machines. Numerical experiments demonstrate its efficacy in solving the TSP, notably solving 8-customer instances, which are among the largest TSPs solved using a QAOA-based approach.
@@
The study by Mohanty at al. mohanty2023analysis presents a VRP solver for three and four cities using VQE on a fixed ansatz. The analysis extends to assess the solution’s robustness in noisy quantum channels, considering various noise models. The work contributes to understanding the impact of noise on quantum algorithms for solving real-world optimisation problems. The results highlight the importance of noise analysis for effective use of quantum devices, emphasising the role of optimisers and the need for further experimentation on physical quantum devices with larger VRP instances.
@@
Bentley et al. bentley2022quantum also introduce a QAOA based framework for integrating quantum computing with the transport sector. They focus on a CVRP and demonstrate a significant circuit performance enhancement (20 times error reduction) using a real quantum device, emphasising problem decomposition, mapping, and compilation for quantum advantage.
@@
Next, Xie et al. xie2023feasibility also use a QAOA approach for the CVRP. This work explores the application of a variant of QAOA called Grover-Mixer Quantum Alternating Operator Ansatz to solve CVRP. The paper introduces a novel binary encoding for CVRP, focusing on minimising the shortest path while bypassing vehicle capacity constraints. The search space is further constrained using the Grover-Mixer. Experimental results indicate that the proposed method outperforms conventional QAOA in terms of feasibility ratios, optimality ratios, and optimality gaps. Despite promising results, the paper acknowledges drawbacks, such as the requirement for multi-controlled Toffoli gates, which may pose challenges for implementation on near-term quantum devices with noise.
@@
Also Palackal et al. palackal2023quantum look at the CVRP and its relation to the TSP. To facilitate experiments, the CVRP is reduced to a clustering phase and a set of TSPs. The authors extensively test QAOA and VQE on TSP instances, investigating the influence of hyperparameters like classical optimizer choice and constraint penalisation strength. Results suggest that QAOA often falls short in reaching the energy threshold for feasible TSP solutions, even with extensions like recursive and constraint-preserving mixer QAOA. On the contrary, VQE performs better, reaching the energy threshold and showing improved performance. The study evaluates the challenges of quantum-assisted solutions for real-world optimisation problems and proposes perspectives for overcoming these obstacles. It also introduces a performance metric tailored to the problem, focusing on solution quality from an application perspective.
@@
Last in this category, the paper by Alsaiyari et al. alsaiyari2023variational compare two quantum algorithms, general VQE and QAOA, for solving the VRP. The study assesses the maximum VRP instance size that the current state-of-the-art NISQ devices can handle, indicating that large VRP instances exceed current quantum capabilities. While recognising the potential of quantum technology, the study emphasises the need for advancements in both hardware and software for broader adoption.
@@
The next paradigm for solving routing problems optimisation is a mixture of quantum computing and machine learning. Sanches et al. sanches2022short explore the integration of quantum computing with reinforcement learning, specifically addressing the VRP. Their model introduces quantum circuits in place of classical attention head layers, showcasing competitive performance. While acknowledging it as a proof of concept, the study emphasises the potential symbiosis between quantum computing and machine learning, raising questions about achieving quantum advantage for combinatorial optimisation through (quantum) reinforcement learning agents.
@@
Correll et al. correll2023quantum take a different approach, employing quantum neural networks with reinforcement learning for truck routing logistics. The study uses real-world data, the same as Weinberg et al. weinberg2023supply we saw earlier, from the automotive sector and achieves results comparable to human decision-making. While the quantum versions of these neural networks offer theoretical speedups, the article acknowledges the need to determine how much of this advantage can be practically realised. Thus, it explores practical realisation methods by exploring methods for tuning quantum circuits and workflow performance.
@@
Slightly different and more general versions of the VRP problem are found in the following papers. Ajagekar et al. ajagekar2020quantum ; ajagekar2020quantum2 propose a Hybrid QC-IQFP Parametric Method for solving the Inexact Quadratic Fractional Programming (IQFP) problem associated with the VRP. The hybrid method combines an inexact parametric algorithm and quantum computing techniques, demonstrating efficiency compared to classical Mixed Integer Nonlinear Programming (MINLP) solvers for medium to large-sized VRP instances. The authors highlight the heuristic nature of the method, its potential for improvement with advancements in quantum computing, and its competitive use in solving large-scale IQFP problems.
@@
Masuda et al. masuda85optimisation evaluate a quantum-classical hybrid approach for the Time-Dependent Vehicle Routing Problem with Time Windows (TDVRPTW). They employ a QUBO formulation using Fixstars Amplify Annealing Engine (a classical, not quantum, Ising machine). The hybrid approach shows potential efficiency for small-scale problems, emphasising the need for further development to handle larger and more complex instances.
@@
Also Leonidas et al. leonidasqubit address the Vehicle Routing Problem with Time Windows (VRPTW). They explore the application of a previously introduced tan2021qubit qubit encoding scheme to reduce the number of binary variables. The study utilizes a quantum variational approach on a testbed of VRPTW instances, ranging from 11 to 3964 routes, formulated as QUBO problems based on realistic shipping scenarios. The authors compare their results with standard binary-to-qubit mappings, employing simulators and various quantum hardware platforms, including IBMQ, AWS (Rigetti), and IonQ. The benchmarking is done against the classical solver, Gurobi. Despite the reduction in qubits required, their approach demonstrates the capability to find approximate solutions to the VRPTW comparable to those obtained from quantum algorithms using the full encoding.
@@
Harwood at al. harwood2021formulating introduce various mathematical formulations for VRPTW, suitable QAOA, VQE, and alternating direction method of multipliers (ADMM). The formulations are compared from a quantum computing perspective, considering metrics for evaluating the difficulty in solving the underlying QUBO problems. Simulated quantum devices are employed to demonstrate the relative benefits of different algorithms and their robustness in practical scenarios.
@@
Irie et al. irie2019quantum introduce the idea of a time-table, capacity-qubits, and the concept of state into the QUBO formulation. Time-tables are incorporated to represent the temporal aspect of VRP, and capacity-qubits are introduced to manage the capacity of vehicles dynamically. The concept of state allows for the description of different traveling rules based on the state of the vehicles. The paper includes a graphical view of a VRP instance solution obtained using the D-Wave 2000Q. The results demonstrate the effectiveness of the proposed formulation for small-size QUBO systems.

### v2

@@ 2.5 Quantum Approximate Optimisation Algorithm
There exists an algorithm that translates QA to the gate-based quantum computer: Quantum Approximate Optimisation Algorithm (QAOA) [47], which is a type of variational quantum eigensolver (VQE) algorithm. The VQE is a quantum algorithm that is used often for quantum chemistry, quantum simulations and optimisation problems. It is a hybrid algorithm that uses both classical computers and quantum computers to find the ground state of a given physical system. In QAOA, the adiabatic evolution is approximated by a discretised function that yields an approximation of the ground state of the desired Hamiltonian. The goal of the QAOA algorithm is to find the lowest possible upper bound for the ground state. Here too, we start with an equal superposition over all solutions |$\ket{s}$. The algorithm then alternately applies the following two operations, operators, to this state:
$U(H_{0},\beta)=e^{-i\beta H_{0}}\textrm{ and }U(H_{1},\gamma)=e^{-i\gamma H_{1}},$ (7)
@@
|$\ket{\gamma,\beta}=U(H_{0},\beta_{p})U(H_{1},\gamma_{p})…U(H_{0},\beta_{1})U(H_{1},\gamma_{1})\ket{s}.$ (8)
Using a QC, the expected value $F_{p}=\bra{\gamma,\beta}H_{1}\ket{\gamma,\beta}$ can then be calculated, which gives an upper limit for the ground state. If $p\rightarrow\infty$ and given the right choice of angels $\beta_{p}$ and $\gamma_{p}$, this approximation will converge to the optimal solution - the exact ground state of the underlying problem. However, finding the right angles is not trivial and if p must be large for a good approximation, the effectiveness of the algorithm is still an uncertain factor. The QAOA has more freedom than QA in this regard. Another choice for the so-called mixing operator $U(H_{0},\beta)$, is the Quantum Alternating Operator Ansatz [61], which can limit the search space, by, for example, setting hard constraints. The QAOA is an example of a variational or hybrid algorithm. Classical parameters are optimised to create a quantum circuit that solves a problem. This class of algorithms are seen to possess the first practical application to gated quantum computers. After all, it only needs a limited number of qubits and the depth of the circuit (the time factor) is also limited if p is not chosen too large. In addition, variational algorithms are fairly resistant to noise, the limited stability of the qubits.
@@ 3.1 Methodology
For this work we created an overview that is as complete as possible until 2023. We used Google Scholar and Scopus as source and used the index terms ‘quantum algorithm’ and the category names of the logistic problems. Note that there exists also a lot of literature on ‘quantum inspired’ algorithms like quantum genetic algorithms [164], quantum evolutionary algorithm [165] and regular simulated annealing, not using the QUBO formulation, such as [32, 102]. The first QA work we found is from 2013 and 2015, with the works from [34, 37, 155]. The main flow of research starts around 2020, as shown in Fig. 3.
@@ 3.2.1 VRP and TSP quantum solutions
One of the earlier works is by Crispin et al. [37]. They propose an approach to solve the Capacitated Vehicle Routing Problem (CVRP) using a QA algorithm. CVRP is a variant of the VRP where vehicles have limited capacity. The paper discusses the design of a spin encoding scheme for CVRP and an empirical approach for tuning parameters in the QA algorithm. QA is introduced as a metaheuristic that utilises quantum tunnelling in the annealing process. Experiments are conducted on CVRP benchmark instances, and the results are compared with a simulated annealing algorithm. The QA algorithm is found to match or outperform simulated annealing in terms of success rate for various instances. The paper analyses the QA algorithm’s performance and discusses the impact of parameters such as population size, temperature, and acceptance ratio.
@@
Srinivasan et al. [147], introduce a quantum algorithm to solve the TSP. It utilises the quantum phase estimation technique [70], encoding distances between cities as phases. Unitary operators are constructed based on these phases, and the phase estimation algorithm is applied to estimate eigenvalues representing total distances for all possible routes. Quantum search algorithms are then employed to find the minimum distance and corresponding route, providing a quadratic speedup over classical brute force methods. The algorithm is illustrated with an example of TSP involving four cities, and simulations are conducted using IBM’s quantum simulator.
Liu at al. [84] introduce a hybrid quantum-classical approach for solving the TSP using path-slicing strategies and quantum local search. By dividing the TSP into smaller subproblems and solving them with quantum annealing and classical solvers, the method efficiently handles computational challenges, achieving near-optimal solutions with improved resource utilization.
@@
The work by Feld et al. [49] also investigates the CVRP using D-Wave’s quantum annealer. They propose a quantum-classical hybrid solution, incorporating a 2-Phase-Heuristic that divides CVRP into clustering and routing phases. Results indicate that the hybrid method can compete with classical construction and 2-phase heuristics in terms of solution quality. However, challenges arise, particularly in finding the ‘best known solution’ for certain datasets, emphasising the need for larger quantum hardware. The study anticipates future improvements in D-Wave’s technology, expecting increased qubit connectivity and quantity.
Other hybrid approaches can be found in Mario et al. [91], presenting a Hybrid Two Step (H2S) and a Hybrid Three Step (H3S) approach and in Sadashivan et al. [126], that combines quantum annealing for the TSP with classical partitioning. Same holds for Osaba et al. [106] presenting a quantum-classical strategy, Q4RPD, for solving real-world package delivery routing problems.
@@
Borowski et al. [19] also delve into (C)VRP, introducing four QA-based algorithms. Their Full QUBO Solver (FQS), Average Partition Solver (AVS), DBSCAN Solver (DBSS), and Solution Partitioning Solver (SPS) showcase promising results. The study employs D-Wave’s Leap framework, comparing quantum algorithms with classical metaheuristics. The experiments use established benchmark datasets and custom scenarios with realistic road networks. The authors outline future research directions, including extending quantum formulations to address more realistic variants of VRP.
@@
Weinberg et al. [159] focus on practical quantum algorithm use, considering the limitations of Noisy Intermediate-Scale Quantum (NISQ) hardware. They adopt a hybrid workflow for solving the CVRP, breaking down problems into smaller instances suitable for QA. Simulated annealing and the D-Wave Hybrid solver are used as quantum substitutes. While not claiming clear advantages over classical methods, they emphasise the sensible application of quantum algorithms to specific bottlenecks, anticipating future advancements.
Kanai et al. [73] explore a hybrid algorithm combining column generation and annealing to solve combinatorial optimization problems with inequality constraints, specifically the CVRP. The proposed method aims to overcome hardware limitations and achieve better lower bounds through annealing-assisted column generation.
@@
This study by Sinno et al. [144] focus on the performance of commercial QA solvers for the CVRP. The research challenges assumptions made in theoretical studies and simulations on classic hardware, emphasising the need for empirical measurements on real quantum platforms to assess commercial capabilities accurately. The study evaluates the quality of solutions provided by the D-Wave CQM solver for CVRP, considering problem size and complexity. The study suggests that, more than problem size, model complexity significantly affects solution quality, highlighting the importance of minimising constraint density in practical applications.
@@
Sales et al. [129] address the VRP in last-mile logistics, proposing a hybrid solution combining QA and classical techniques. Their analysis, using Amazon Braket and D-Wave’s QA, demonstrates the superiority of the hybrid algorithm over full QA, suggesting potential advantages in scenarios with lower data “clusterability.” Despite current challenges, the study identifies a cost-effective relationship between running quantum algorithms and result quality, indicating quantum computing’s theoretical advantage in solving constrained clustering problems.
@@
The thesis by Palmieri [109] explores the application of quantum optimisation techniques to address the (C)VRP. Here, a heuristic two-phase approach is explored. Nodes are first clustered into disjoint groups, and routing is performed individually within each cluster. Various classical methodologies for the clustering phase are investigated. An alternative approach models the clustering phase as a Modularity Maximization problem, which is tackled using quantum algorithms. The routing phase is modelled as a Travelling Salesman Problem (TSP), and subtour elimination constraints are adaptively added using the Dantzig-Fulkerson-Johnson (DFJ) strategy. This adaptive approach reduces the number of qubits, making the models suitable for both gate-based and annealing-based quantum devices. Various quantum algorithms, such as QA, quantum GAVA (Graver Augmented Multi-seed Algorithm), QAOA and VQE are investigated for quantum optimisation. Gurobi is used as a classical benchmarking method. QA and quantum GAMA outperform the VQE approaches here.
@@
Until here, all works use QA as their solver for the VRP. Rana et al. [121] explore employee transport route and agri-tech logistics optimisation using both a quantum gate-based variational algorithm and QA. Here again, quantum hardware limitations lead to problem simplifications, using problem-specific decomposition techniques, such as distance-based clustering. The study utilises real-world data and anticipates extending quantum methods to more complex scenarios, acknowledging ongoing improvements in the D-Wave’s system for addressing larger, intricate logistics optimisation problems.
@@
Azad et al. [12] explore the application of the QAOA for the VRP within Intelligent Transportation Systems (ITS). The paper provides an Ising formulation for the VRP and its variants, adapting them for quantum processing. Despite the limitations of NISQ devices, having 15 qubits, QAOA holds promise for solving combinatorial optimisation problems like VRP. However, they conclude that there is a need for a multi-faceted approach to enhance its performance. Working on aspects such as initialisation techniques, optimising parameter values, exploring different mixer Hamiltonians, and addressing noise-related challenges are essential for realising the full potential of QAOA in practical applications.
@@
The proposed Indirect QAOA (IQAOA) by Bourreau et al. [20], introduces a novel approach, targeting the Hamiltonian of a set of string vectors. It utilises a Quantum Alternating Operator Ansatz with a parameterised family of unitary operators, allowing for efficient modelling of the Hamiltonian. This algorithm addresses the limitations of standard QAOA by incorporating a classical meta-optimisation loop and estimating the average cost of each string vector. The IQAOA approach includes the creation of a quantum circuit with a limited number of gates suitable for current noisy quantum machines. Numerical experiments demonstrate its efficacy in solving the TSP, notably solving 8-customer instances, which are among the largest TSPs solved using a QAOA-based approach.
@@
The study by Mohanty at al. [99] presents a VRP solver for three and four cities using VQE on a fixed ansatz. The analysis extends to assess the solution’s robustness in noisy quantum channels, considering various noise models. The work contributes to understanding the impact of noise on quantum algorithms for solving real-world optimisation problems. The results highlight the importance of noise analysis for effective use of quantum devices, emphasising the role of optimisers and the need for further experimentation on physical quantum devices with larger VRP instances.
@@
Bentley et al. [16] also introduce a QAOA based framework for integrating quantum computing with the transport sector. They focus on a CVRP and demonstrate a significant circuit performance enhancement (20 times error reduction) using a real quantum device, emphasising problem decomposition, mapping, and compilation for quantum advantage. Same holds for the approaches in Ramezani et al. [120] and Herzog et al. [65], that propose circuit cutting and other techniques to reduce the circuit size and the number of required qubits.
@@
Next, Xie et al. [163] also use a QAOA approach for the CVRP. This work explores the application of a variant of QAOA called Grover-Mixer Quantum Alternating Operator Ansatz to solve CVRP. The paper introduces a novel binary encoding for CVRP, focusing on minimising the shortest path while bypassing vehicle capacity constraints. The search space is further constrained using the Grover-Mixer. Experimental results indicate that the proposed method outperforms conventional QAOA in terms of feasibility ratios, optimality ratios, and optimality gaps. Despite promising results, the paper acknowledges drawbacks, such as the requirement for multi-controlled Toffoli gates, which may pose challenges for implementation on near-term quantum devices with noise.
Also Sato et al. [133] use Grover’s algorithm for the initial state preparation in quantum circuits, followed by amplification of the optimal solution. Their method leverages Higher-Order Unconstrained Binary Optimization (HOBO) encoding to reduce qubit requirements and improve efficiency.
@@
Also Palackal et al. [108] look at the CVRP and its relation to the TSP. To facilitate experiments, the CVRP is reduced to a clustering phase and a set of TSPs. The authors extensively test QAOA and VQE on TSP instances, investigating the influence of hyperparameters like classical optimizer choice and constraint penalisation strength. Results suggest that QAOA often falls short in reaching the energy threshold for feasible TSP solutions, even with extensions like recursive and constraint-preserving mixer QAOA. On the contrary, VQE performs better, reaching the energy threshold and showing improved performance. The study evaluates the challenges of quantum-assisted solutions for real-world optimisation problems and proposes perspectives for overcoming these obstacles. It also introduces a performance metric tailored to the problem, focusing on solution quality from an application perspective.
Li et al. [83] explore variational quantum algorithms for collision-avoidance route planning in consumer electronics supply chains. They introduce infeasible solution constraints to optimize qubit usage, integrates vehicle routing with collision-avoidance, and validates the method through quantum simulations, enhancing scalability with a stepwise optimization algorithm.
@@
Last in this category, the paper by Alsaiyari et al. [6] compare two quantum algorithms, general VQE and QAOA, for solving the VRP. The study assesses the maximum VRP instance size that the current state-of-the-art NISQ devices can handle, indicating that large VRP instances exceed current quantum capabilities. While recognising the potential of quantum technology, the study emphasises the need for advancements in both hardware and software for broader adoption.
@@
The next paradigm for solving routing problems optimisation is a mixture of quantum computing and machine learning. Sanches et al. [132] explore the integration of quantum computing with reinforcement learning, specifically addressing the VRP. Their model introduces quantum circuits in place of classical attention head layers, showcasing competitive performance. While acknowledging it as a proof of concept, the study emphasises the potential symbiosis between quantum computing and machine learning, raising questions about achieving quantum advantage for combinatorial optimisation through (quantum) reinforcement learning agents.
@@
Correll et al. [36] take a different approach, employing quantum neural networks with reinforcement learning for truck routing logistics. The study uses real-world data, the same as Weinberg et al. [159] we saw earlier, from the automotive sector and achieves results comparable to human decision-making. While the quantum versions of these neural networks offer theoretical speedups, the article acknowledges the need to determine how much of this advantage can be practically realised. Thus, it explores practical realisation methods by exploring methods for tuning quantum circuits and workflow performance.
The paper by Xu et. al [166] proposes a quantum Q-learning model for the CVRP. It uses parameterised quantum circuits to approximate Q-values, enhancing solution quality and efficiency. Experimental results show that the model outperforms classical methods and previous quantum approaches, making it more suitable for NISQ devices.
Also quantum support vector machines are used to solve VRP problems, by Mohanty et al. [100]. The combine this with variational quantum eigensolvers on 6- and 12-qubit circuits for 3- and 4-city scenarios. Various quantum encoding techniques and optimizers are evaluated using IBM Qiskit simulations.
Lastly we see the rise of quantum variants of the metaphor based meta-heuristics. Qiu et al. [118] introduce a hybrid quantum-classical algorithm that enhances Quantum Ant Colony Optimization (QACO) by integrating a clustering algorithm, specifically K-means. The proposed method was tested using the TSP and demonstrated improved performance across multiple datasets. Additionally, the hybrid algorithm showed robustness to noise, a critical challenge in quantum computing. The study highlights the potential of QACO in the Noisy Intermediate-Scale Quantum (NISQ) era by extending its applicability through classical-quantum integration.
@@
Slightly different and more general versions of the VRP problem are found in the following papers. Ajagekar et al. [3, 4] propose a Hybrid QC-IQFP Parametric Method for solving the Inexact Quadratic Fractional Programming (IQFP) problem associated with the VRP. The hybrid method combines an inexact parametric algorithm and quantum computing techniques, demonstrating efficiency compared to classical Mixed Integer Nonlinear Programming (MINLP) solvers for medium to large-sized VRP instances. The authors highlight the heuristic nature of the method, its potential for improvement with advancements in quantum computing, and its competitive use in solving large-scale IQFP problems.
@@
Masuda et al. [95] evaluate a quantum-classical hybrid approach for the Time-Dependent Vehicle Routing Problem with Time Windows (TDVRPTW). They employ a QUBO formulation using Fixstars Amplify Annealing Engine (a classical, not quantum, Ising machine). The hybrid approach shows potential efficiency for small-scale problems, emphasising the need for further development to handle larger and more complex instances.
@@
Also Leonidas et al. [82] address the Vehicle Routing Problem with Time Windows (VRPTW). They explore the application of a previously introduced [150] qubit encoding scheme to reduce the number of binary variables. The study utilizes a quantum variational approach on a testbed of VRPTW instances, ranging from 11 to 3964 routes, formulated as QUBO problems based on realistic shipping scenarios. The authors compare their results with standard binary-to-qubit mappings, employing simulators and various quantum hardware platforms, including IBMQ, AWS (Rigetti), and IonQ. The benchmarking is done against the classical solver, Gurobi. Despite the reduction in qubits required, their approach demonstrates the capability to find approximate solutions to the VRPTW comparable to those obtained from quantum algorithms using the full encoding.
@@
Harwood at al. [63] introduce various mathematical formulations for VRPTW, suitable QAOA, VQE, and alternating direction method of multipliers (ADMM). The formulations are compared from a quantum computing perspective, considering metrics for evaluating the difficulty in solving the underlying QUBO problems. Simulated quantum devices are employed to demonstrate the relative benefits of different algorithms and their robustness in practical scenarios.
@@
Irie et al. [67] introduce the idea of a time-table, capacity-qubits, and the concept of state into the QUBO formulation. Time-tables are incorporated to represent the temporal aspect of VRP, and capacity-qubits are introduced to manage the capacity of vehicles dynamically. The concept of state allows for the description of different travelling rules based on the state of the vehicles. The paper includes a graphical view of a VRP instance solution obtained using the D-Wave 2000Q. The results demonstrate the effectiveness of the proposed formulation for small-size QUBO systems.